OpenStudy (ayanaortega):

can someone help me please factor completely. 3x^2+21x+30 @563blackghost

OpenStudy (ayanaortega):

show work pleasl

OpenStudy (freckles):

do you see that 3,21,30 all share a common factor?

OpenStudy (ayanaortega):

yes

OpenStudy (freckles):

and what common factor is that?

OpenStudy (ayanaortega):

3

OpenStudy (freckles):

very awesome can you tell me what you get after dividing the following numbers 3,21, and 30 by 3?

OpenStudy (ayanaortega):

1

OpenStudy (freckles):

3/3=1 21/3=? 30/3=?

OpenStudy (ayanaortega):

21/3=7 30/3=10

OpenStudy (freckles):

so you should see that we have the following so far \[3x^2+21x+30 \\ 3(1x^2+7x+10) \\ 3(x^2+7x+10)\] (we are not done)

OpenStudy (freckles):

does this seem okay so far?

OpenStudy (ayanaortega):

yes

OpenStudy (freckles):

okay now let's look at just the x^2+7x+10 for now

OpenStudy (freckles):

we will bring the 3 down later

OpenStudy (freckles):

so looking at \[x^2+7x+10\] we need to come up with 2 numbers that multiply to be 10 while also adds up to be 7

OpenStudy (freckles):

if the coefficient of x^2 was something other that 1 say a I would have said we need to come up with 2 numbers that multiply to be a*10 while adds up to be 7 that is if we had something like ax^2+7x+10 where a is given as some number

OpenStudy (freckles):

can you think of two numbers that multiply to be 10 while they also add up to be 7?

OpenStudy (ayanaortega):

2 and 5

OpenStudy (ayanaortega):

@freckles

OpenStudy (freckles):

right 2*5 is 10 while 2+5=7 so you can do this the short cut way since the coefficient of x^2 is 1 which is if you have x^2+bx+c and you find two numbers m and n such that m*n=c while m+n=b then you can say x^2+bx+c is the same as (x+m)(x+n) since when multiplying this out you get x^2+mx+nx+mn =x^2+(m+n)x+mn=x^2+bx+c so all you need to do when you find m and n such that m*n=b while m+n=c is write x^2+bx+c as (x+m)(x+n) or if you want to do it the longer way the way that works for if you have something other than 1 as the coefficient of x^2 then you would do something like this ax^2+bx+c look for two numbers that multiply to be a*c and add up to b say that numbers are m and n so you replace bx with mx+nx ax^2+mx+nx+c and factor by grouping

OpenStudy (freckles):

if you don't feel like reading all of that right now then just read this part... " so all you need to do when you find m and n such that m*n=b while m+n=c is write x^2+bx+c as (x+m)(x+n) "

OpenStudy (freckles):

so using what I just said rewrite x^2+7x+10

OpenStudy (ayanaortega):

okay i rewright

OpenStudy (freckles):

and remember your m and n are 2 an 5 correct?

OpenStudy (ayanaortega):

yes

OpenStudy (freckles):

so you know you are going to write x^2+7x+10 as (x+ m )(x+ n ) where m and n are 2 and 5 it doesn't matter if you replace m with 5 or 2 just the vice versa for n

OpenStudy (ayanaortega):

this is hard this is i have so far 3x^2+21x+30 3(1x^2+7x+10) 3(x^2+7x+10) x^2+7x+10 (x+m)(x+n)

OpenStudy (freckles):

you found your values for m and n already....

OpenStudy (ayanaortega):

no thats the hard part

OpenStudy (freckles):

x^2+bx+c in factored form is (x+m)(x+n) if m*n is c and m+n is b x^2+7x+10 in factored form is ..... since 2*5=10 and 2+5=7 what goes in place of the dots

OpenStudy (ayanaortega):

X

OpenStudy (freckles):

no x^2+7x+10 is not X

OpenStudy (freckles):

you already found m and n you just need to replace m and n

OpenStudy (freckles):

remember you said 2*5=10 and 2+5=7 and in my general thing I have above m*n=c and m+n=b and my factored form for x^2+bx+c was (x+m)(x+n)

OpenStudy (freckles):

you found m and n to be 2 and 5

OpenStudy (freckles):

you just need to replace m and n now in the factored form for the quadratic I wrote above

OpenStudy (ayanaortega):

1x^2

OpenStudy (freckles):

are you reading what I wrote?

OpenStudy (ayanaortega):

yes

OpenStudy (freckles):

The factored form for x^2+bx+c is (x+m)(x+n) where m*n=c and m+n=b You have the factored form for x^2+7x+10 is ....... where 2*5=10 and 2+5=7 just replace the m and n in my factored form for the quadratic with your m and n which is 2 and 5 and you will have your factored form for your quadratic

OpenStudy (ayanaortega):

so 2 and 5 is the replacement right

OpenStudy (freckles):

example the factored form for x^2+5x+6 is (x+2)(x+3) since 2*3=6 and 2+3=5 another example the factored form for x^2-5x+6 is (x-2)(x-3) since -2(-3)=6 and -2+(-3)=-5

OpenStudy (freckles):

another example the factored form for x^2-x-6 is (x-3)(x+2) since -3(2)=6 and -3+2=-1

OpenStudy (freckles):

in general The factored form for x^2+bx+c is (x+m)(x+n) if m*n=c and m+n=b

OpenStudy (freckles):

your problem: the factored form for x^2+7x+10 is .................... since 2*5=10 and 2+5=7

OpenStudy (freckles):

another example the factored form for x^2+x-6 is (x+3)(x-2) since 3(-2)=6 and 3+(-2)=1 another example the factored form for x^2+6x+8 is (x+2)(x+4) since 2(4)=8 and 2+4=6

OpenStudy (ayanaortega):

so 8 and 6

OpenStudy (ayanaortega):

wait no 4

OpenStudy (freckles):

what are you saying about my example?

OpenStudy (ayanaortega):

7x and 10 is the replacement

OpenStudy (freckles):

I don't understand why you are having trouble with this... the factored form for x^2+7x+10 is (x+2)(x+5) since 2*5=10 and 2+5=7

OpenStudy (ayanaortega):

so 2and 5 is the replacement for m and n right

OpenStudy (freckles):

yes this has been mentioned before :p

OpenStudy (freckles):

x^2+bx+c has formed form (x+m)(x+n) if m*n=c and m+n=b since we had 2*5=10 and 2+5=7 then m=2 and n=5 or vice versa m=5 and n=2 doesn't matter

OpenStudy (freckles):

anyways I have to go for now I have mentioned several examples above of using the general thing I mentioned which was " The factored form for x^2+bx+c is (x+m)(x+n) if m*n=c and m+n=b "